OpenStudy (anonymous):

imagine two solutions with the same concentration and the same boiling point but one has ethanol as the solvent and the other has carbon tetrachloride

OpenStudy (anonymous):

Someone have asked the same question before. :) I do not take credit boiling point = 80.1 - 76.8 = 3.3 2.53m + 3.3 = 5.03m 2.53m - 5.03m = -3.3 ===> m = [1.32] dT = kb*m dT = 2.53 * 1.32 = 3.34 Tb=> 80.1 + 3.34 = [83.44] or dT = 5.03 * 1.32 = 6.63 Tb=> 76.8 + 6.63 = [83.44] same answer with benzene